OpenStudy (anonymous):

Simplify: -3(5x^3-3x^2-9x-1)-4(-2x^3-8x^2+12) =

OpenStudy (anonymous):

-7x^3+41x^2+27-45 ?????????????????????????????????

OpenStudy (psymon):

For a problem like this, it's almost more about being careful than anything else. I think it's good to get into a habit of not dropping parenthesis and maybe even trying to add them to make sure you get signs and simplification correct. But anyway... I'll start by distributing in the -3 into the first group of parenthesis. I'll also keep those parenthesis in tact JUST in case. I'll list each multiplication result in a list to help me keep track -3*5x^3 = -15x^3 -3*-3x^2 = 9x^2 -3*-9x = 27x -3*-1 = 3 So my first group is (-15x^3 + 9x^2 + 27x + 3) Now on to the second group. That minus needs to distribute into the 2nd group as well as the 4. So this is the list of each multiplication -4*-2x^3 = 8x^3 -4*-8x^2 = 32x^2 -4*12 = -48 So my second group is +(8x^3 + 32x^2 -48). So now I'll just group like terms. -15x^3 + 8x^3 = -7x^3 9x^2 + 32x^2 = 41x^2 27x 3 - 48 = -45 Combined together I finish with -7x^3 + 41x^2 + 27x - 45 I know you messaged me, but I just wanted to finish the explanation. You can message other questions if you want afterwards, I have no issue with that.

OpenStudy (anonymous):

thank you very much!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (psymon):

Of course :3